OpenStudy (anonymous):

Medal and follow ...... If the nth term of this sequence (6 9 12 15 18) is represented by f(n) which of the following functions best represents this sequence?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@math

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

Help me

OpenStudy (anonymous):

@undeadknight26

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

Can anyone help

OpenStudy (campbell_st):

ok... so the 1st term is 3 and it increases by 3... so the general term is \[a_{n} = 3 + (n - 1)\times 3\] just simplify it

OpenStudy (anonymous):

Is it f(n) = 3n + 3; n>= 1 Or f(n) = 3n - 3; n>= 1 Or nethier

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

well what happens with f(n) = 3 +(n - 1) 3 f(n) = 3+ 3n - 3 you need to simplify this...

OpenStudy (anonymous):

Ohhhhhhh i get it 6n -3 n is equal to oupr greater than 1